OpenStudy (anonymous):

the key to using the elimantion method is to find variable terms in the two equations that have unequal coefficients true or false?

OpenStudy (anonymous):

would opposite coefficients be the same as unequal coefficients ? For example : you would want 1 equation to have 4x and the other equation to have -4x, so that they can cancel each other out.

OpenStudy (anonymous):

@amistre64 ....question for you ?

OpenStudy (anonymous):

@whpalmer4 .....I have a question

OpenStudy (whpalmer4):

you've got questions, we've got answers :-)

OpenStudy (whpalmer4):

but first, often more questions to you :-)

OpenStudy (anonymous):

are opposite coefficients the same as unequal coefficients ?

OpenStudy (anonymous):

I think they are, but I want to confirm

OpenStudy (whpalmer4):

well, I'm not sure what whoever is writing this considers "opposite". the way I say it is that the magnitude needs to be equal, and the sign opposite. so, 4 and -4 is a good matchup, because the magnitude of both is 4, and the sign is opposite. 4 and -3 is not a good matchup for doing elimination, because although the signs are opposite, the magnitude is not the same.

OpenStudy (whpalmer4):

really, the key is that the magnitudes are equal. it doesn't matter if the signs are identical or opposite because you can either subtract or add to combine the equations.

OpenStudy (whpalmer4):

it is perhaps easier if the signs are opposite, just because many people make fewer mistakes adding than they do subtracting, especially if negative numbers are involved.

OpenStudy (anonymous):

so would you consider the answer to this question true or false ?

OpenStudy (anonymous):

they are unequal

OpenStudy (anonymous):

-4 is not equal to 4...right ?

OpenStudy (anonymous):

although they are both 4 spaces from 0

OpenStudy (anonymous):

I am just confused on the wording

OpenStudy (whpalmer4):

I would consider it false, because you can eliminate with equal coefficients. Let me demonstrate: \[2x + 3y = 5\]\[-2x + y = 7\]I can eliminate here by simply adding the two together: \[(2-2)x + (3+1)y = 5+7\]\[4y=12\] but what if I had \[2x + 3y =5 \]\[2x -y = -7\] (the second equation is equivalent to the second equation of the first set, just multiplied through by -1) Well, here I can subtract! \[2x + 3y = 5\]\[2x-y=-7\]------------ \[(2-2)x + (3-(-1))y = 5-(-7)\]\[4y = 12\] same result!

OpenStudy (whpalmer4):

so, the key is that the magnitude was the same, not that the signs were opposite.

OpenStudy (whpalmer4):

magnitude is just another word for absolute value in the case of a single number...

OpenStudy (anonymous):

thanks because I was leaning towards true......I see the light...lol....

OpenStudy (whpalmer4):

I prefer to do it by addition so I don't risk making mistakes doing subtraction of negative numbers. If I have to multiply one equation by -1 to make the signs opposite, I do so, that's easy.

OpenStudy (amistre64):

the key to using the elimantion method is to find variable terms in the two equations that have unequal coefficients not necessarily

OpenStudy (whpalmer4):

personally, I'm not a fan of questions like this that perhaps are a better test of your ability to parse the question than your understanding of the underlying concept...

OpenStudy (amistre64):

x + 3y = 6 x - 2y = 7 the xs have the same coeff, but its just as easily eliminated

OpenStudy (anonymous):

thats true....multiplying by -1.....I understand now....the wording threw me off

OpenStudy (whpalmer4):

not that parsing questions accurately isn't a valuable skill to develop, but presumably this is a math class, not a critical reading class :-)

OpenStudy (anonymous):

thanks for explaining

OpenStudy (whpalmer4):

you're welcome!

OpenStudy (anonymous):

thank you everyone this really helped and i learned a few things too ^_^